# Positional encoding is not the same as context: A study on positional encoding for Sequential recommendation

### v1

@@ 3.1 Positional encodings
enables models to weigh different parts of input sequences differently, allowing them to focus on relevant information during processing. The self-attention mechanism was introduced in Vaswani et al. (2017), which laid the foundation for the transformer architecture.
In self-attention, the input sequence is typically represented as a set of vectors, where each vector corresponds to a position in the sequence. This allows each position to attend to all other positions and captures dependencies and relationships within the sequence. For a given dimensionality $d$ of the item representation and a hidden dimension $H<d$, the attention weight assigned to each position is computed based on the similarity (dot product) between the query $Q\in\mathbb{R}^{d\times H}$, key $K\in\mathbb{R}^{d\times H}$, and value $V\in\mathbb{R}^{d\times H}$ matrices associated with the positions. These learnable matrices are derived from linear transformations of the original input $X\in\mathbb{R}^{n\times d}$ by multiplication, where $n$ is the number of items in the sequence.
The attention weight ($\alpha_{ij}$) assigned to position $i$ for attending to position $j$ is computed using the following formula:
$\alpha_{ij}=\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K)^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{t}\,K)^{T}/\sqrt{d})}}$ (1)
Here, $Q_{i}$ and $K_{j}$ represent the query and key matrices associated with positions $i$ and $j$, respectively. The dot product is scaled by $\sqrt{d}$ to prevent vanishing gradient problems. The resulting attention weights are then used to compute a weighted sum of the value matrices ($V_{j}$) associated with all positions:
$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}X\cdot V_{j}$ (2)
This weighted sum, representing the attended information, is then used as the output of the self-attention mechanism. The process is repeated for all positions in the sequence.
Absolute Positional Encoding (APE )
Absolute positional encoding adds fixed values to the embeddings of items based on their absolute position in the sequence. This means all the items have the same initial token as a reference. The most common approach is based on trigonometry, like in BERT Devlin et al. (2019), the so-called sinusoidal encoding. Absolute encoding is added at the input layer as part of the item encoding, 1. The most common practice is to add a vector encoding the position.
The absolute positional encoding for a token at position $i$ in a sequence of length $N$ can be calculated using trigonometric functions such as sine and cosine. A common formula is:
$PE_{(pos,2i)}=\sin\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (3)
$PE_{(pos,2i+1)}=\cos\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (4)
where $pos$ is the position of the item, $i$ is the dimension, and $d$ is the dimension of the embedding. This encoding is usually added to the original embedding, and we will refer to it along the text as absolute encoding (AE). For a given item embedding $x$ at position $pos$, we replace it as follows:
$x=x+PE_{(pos,*)}$ (5)
Another way of making the positional information more explicit and separated from the rest of the context and attributes is by concatenation, where we divide the dimension of the embedding into the context and attribute information and the positional encoding information:
$x=Concat\,(x,PE_{(pos,*)})$ (6)
We apply a Linear layer right after to recover the original dimension and do not change the number of parameters in the model. We will refer to them as concat absolute encoding (CAE). We added the position information orthogonally to make the model gradually rely on the position, producing more stable results, Section 5.
Learnable Encoding
Learnable position embeddings take a different approach by treating the positional information as parameters that can be updated during training. Instead of relying on predefined patterns, the model learns the optimal representation of positions based on the data context. We will denote these embeddings as learnable encoding (LE). In this case, we add a positional encoding, which is merely a learnable vector. We also apply a concatenation approach denoting as concatenated learnable encoding (CLE). In this case, their results will be generally worse than the classic residual connection.
Relative Positional Encoding (RPE )
A disadvantage of using APE is that it requires a fixed length of the input sequence and does not directly capture the relative positions of one item to another, whether the other is further ahead or backward. Relative position encoding (RPE ) was proposed to solve these issues, 1. Shaw et al. (2018) uses relative position encoding (RPE ) instead of APE and adds the position embeddings to the key and optional value projections instead of the input. A simplified version deleting the position embeddings in value projections in TranformersXL Dai et al. (2019) showed better performance on language modelling tasks. The length of the user’s session or action log varies significantly from user to user, suggesting that these encoding techniques may be of value to SRS.
The relative positional encoding can be computed using trainable parameters. The specific formula may depend on the implementation; in our case, we follow the one in Shaw et al. (2018), where the position is given inside the attention. The relation between two inputs $X_{i}$ and $X_{j}$ is given by $a_{ij}^{V},a_{ij}^{K}\in\mathbb{R}^{d_{a}}$, for values and keys. This information can be shared among heads, but we consider the case in which information is not shared.
For the relative case, the attention weight ($\alpha_{ij}$) can be written as:
$\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}$ (7)
We also introduce the position at the value product as follows:
$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}(X\cdot V_{j}+a_{ij}^{V})$ (8)
Relative positional encoding can be especially useful when dealing with long sequences, as it allows the model to generalise better to different distances between items and in cases where the local position information is more critical. As a drawback, since it has to be applied to each attention head, it tends to be slower Shaw et al. (2018). We will denote these embeddings as RMHA-4 , where we bound the relative indexes to a distance of $4$. We do not apply a concatenated version in this case.
Rotary Encoding (Rope and Rotatory )
Unlike absolute or relative positional encoding, which focuses on representing positions in a fixed coordinate system or relative to other items in a sequence, rotary encoding leverages rotational transformations to encode position information. "Rotary Positional Encoding" typically refers to a specific technique used in positional encoding, Su et al. (2024). We will use the notation Rope to refer to them.
The motivation behind Rope is that the orientation or rotation of positional embeddings can provide valuable information about the relationship between items in a sequence. This approach is particularly relevant in scenarios where the inherent order or arrangement of items is not straightforward and can benefit from a more flexible representation, like a user searching for more than one item in the same session. These embeddings capture both linear order and rotational relationships within the sequence. Rotary position encodings apply a rotation to the Query and Key matrices, which preserves the dot product among the vectors.
In the two dimensional case for a given input vector $x$ the encoding function $f$ for given angle $\theta$ is defined as follows:
$f_{\theta}(\mathbf{x})=\mathbf{R}(\theta)\cdot\mathbf{W}\cdot\mathbf{x}$ (9)
Here, $\mathbf{R}(\theta)$ represents the rotation matrix parameterized by an angle $\theta$, expressed using sines and cosines, and $\mathbf{W}$ is either $K$ or $Q$. More explicit, for the $K$ case we have:
$\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\cdot\begin{bmatrix}K_{11}&K_{12}\\ K_{21}&K_{22}\end{bmatrix}\cdot\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$ (10)
The straightforward generalisation of this to higher dimensionality divides the vector dimension into two-dimensional chunks and applies the rotation to each of them. Since this implementation is costly, the actual implementation expresses this rotation as two vector multiplications and one vector addition, as RoFormer Su et al. (2024).
Since we want to analyse the effect of the rotational position for sinusoidal, we propose here a new encoding, Rotatory , which is concatenated version Rotatory + Con. . Instead of directly applying sinusoidal functions, rotary encodings introduce an additional rotation mechanism by separating the angles into sine and cosine components and then alternating them across the embedding dimension. By alternating sine and cosine embeddings, the model can capture different phases of positional information, effectively introducing a rotational component to the encoding.
This rotation helps mitigate potential issues, such as vanishing gradients that can arise with traditional positional encoding, which is especially important for stability. So, while the term "rotary" might not directly imply physical rotation, it refers to the rotational mechanism introduced in the encoding process to enhance the representation of positional information in Transformer models. The implementation is similar to that from Rope , but applying the rotation matrices to the vector encodings added to the input items.
Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
@@
Beauty silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97
Beauty leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48
@@
Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
@@
Table 2: The table displays the top $4$ results for each dataset on cases with a standard deviation for $HIT$ below $3$.
Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
@@
Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
@@
Table 3: he table displays the top $4$ results for each dataset on cases with a standard deviation for HIT below $12$.
4 Experimental Setup
4.1 Datasets
We selected the Amazon dataset to assess the proficiency of various encoding types. For this purpose, we utilised four different real-world datasets extracted from product reviews on Amazon.com. These datasets have been diverse and widely used for SRS under the leave-one-out protocol (He and McAuley (2016b), Kang and McAuley (2018b), Hou et al. (2019), Wu et al. (2020), Steck (2019) and Zhou et al. (2020)). Since multiple versions of these datasets exist, we opted for the preprocessed versions from Rashed et al. (2022).
These datasets encompass various product categories, providing a diverse and comprehensive source for training and evaluating recommendation models. They provide essential features for each interaction, such as user IDs, item IDs, timestamps, and additional vector contextual information. As we can see in Tables 5, and 4, they provide diversity in the sparsity, which is going to be a key point in our analysis, but also on the discreteness of the features.
Dataset Users Items Interactions Attributes
Men 34,244 110,636 254,870 2,048
Fashion 45,184 166,270 358,003 2,048
Games 31,013 23,715 287,107 506
Beauty 52,204 57,289 394,908 6,507
Table 4: Dataset Statistics.
1.Beauty
(Wu et al. (2020), Zhou et al. (2020) and Kang and McAuley (2018b)) The dataset includes discrete and categorical attributes for all beauty products, including fine-grained categories and brands. Mainly categorical and discrete features.
2.Video Games
(Wu et al. (2020) and Kang and McAuley (2018b)) The Video Games sub-dataset includes user interactions, reviews, and product details specific to the video game category as price, brand and categorical features. Most of the attributes here are discrete and categorical.
3.Men
(He and McAuley (2016b)) The men’s dataset encompasses a comprehensive collection of items falling under men’s clothing. The attributes are dense vectors from image-based features extracted from the last layer of a ResNet50 Deng et al. (2009) on the ImageNet dataset He et al. (2016a).
3.Fashion
(He and McAuley (2016b) and Hou et al. (2019)) It contains 6 categories for men’s and women’s clothing. The features were extracted as dense using the same ResNet50 approach.
4.2 Models and Training
To implement CARCA, we converted TensorFlow code from the original model into Pytorch [^1] https://github.com/researcher1741/Position_encoding_SRS. We extracted the implementation in Papso (2023) for the part of the loss. The original version in TensorFlow was written on top of previous models written in TensorFlow. We found that these implementations check the metrics every $20$ epoch, assuming that the overfitting starts later. This is probably because some models need more than $1000$ epochs, and the cost of measurement becomes significant. We measure the metrics in a logarithmic manner, giving rise to slightly better results in some cases. Our implementation can be found here .
As in previous works (Rashed et al. (2022), Kang and McAuley (2018a), etc.), we use an ADAM optimiser to minimise the binary cross-entropy loss of the CARCA model while masking the padded items to prevent them from contributing to the loss function. For a given sequence of items interactions for a user $u$ as $\{i^{u}_{1},\ldots,i^{u}_{N}\}$, we create an input list as $I^{u+}=\{i^{u}_{1},\ldots,i^{u}_{N-1}\}$ by deleting the last item of the list, a positive target list as $T^{u+}=\{i^{u}_{2},\ldots,i^{u}_{N}\}$ by shifting one the input list, and a negative target list as $T^{u-}=\{i^{rand_{1}},\ldots,i^{rand_{N}-1}\}$ generated by random negative items. Then the loss is given by:
4.3 Metrics
We use two commonly used metrics for Next-Item Recommendations/SRS. Hit Rate measures the proportion of recommended items that overlap with the user’s actual interactions. Hit@10 counts the fraction of times the ground truth next item is among the top 10 items.

### v2

@@ 2.2. Positional encodings
This mechanism allows models to weigh different parts of the input sequence differently, focusing on relevant information during processing. Introduced in (Vaswani et al., 2017), the self-attention mechanism computes attention weights based on the similarity between queries ($Q$), keys ($K$), and values ($V$):
(4) $\alpha_{ij}=\frac{\exp\left((Q_{i}K_{j}^{T})/\sqrt{d}\right)}{\sum_{t=1}^{n}\exp\left((Q_{i}K_{t}^{T})/\sqrt{d}\right)}$
(5) $\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}V_{j}$
Relative Positional Encoding Implementation (RMHA-4 )
This encoding addresses the limitations of APE by encoding the relative positions within the attention mechanism. Models like NARM (Li et al., 2017) and TransformerXL (Dai et al., 2019) adopt RPE to enhance the modeling of contextual dependencies. In RPE , positional information is incorporated into the key and value projections, allowing the model to focus on the relationships between items dynamically.
The attention weight with RPE is computed as:
(6) $\alpha_{ij}=\frac{\exp\left((Q_{i}(K_{j}+a_{ij}^{K})^{T})/\sqrt{d}\right)}{\sum_{t=1}^{n}\exp\left((Q_{i}(K_{t}+a_{it}^{K})^{T})/\sqrt{d}\right)}$
Additionally, positional information can be added to the value projections:
(7) $\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}(V_{j}+a_{ij}^{V})$
We will consider the case which bounds the relative indexes to a distance of $4$ and called it RMHA-4 , Relative Multi-Head Attention with distance $4$.
RoPE (RoPE )
Rotary Positional Encoding (Su et al., 2024) encodes positional information by applying rotational transformations to the Query and Key matrices. This preserves the dot product between vectors, maintaining effective attention mechanisms. RoPE is particularly useful for complex item orders, providing a more flexible representation than traditional encodings.
For a two-dimensional vector $\mathbf{x}$ and angle $\theta$:
(8) $f_{\theta}(\mathbf{x})=\mathbf{R}(\theta)\mathbf{W}\mathbf{x}$
where $\mathbf{R}(\theta)$ is a rotation matrix defined by sines and cosines, and $\mathbf{W}$ represents either $K$ or $Q$. Higher-dimensional vectors are split into two-dimensional chunks, with rotations applied to each. To reduce computational costs, RoFormer (Su et al., 2024) implements these rotations using two vector multiplications and one vector addition.
3. Methodology
In this section, we present our contributions to positional encodings by introducing variants of existing methods and new encodings.
3.1. Encoding Variants
The most commonly used encodings are Abs encodings, which are vector encodings added to the initial item vector by addition. This approach preserves the original embedding dimension without requiring additional encoding dimensions (Table 1). In contrast, non-vector encodings, such as Relative Positional Encoding (RPE) and Rotary Positional Encoding (RoPE), are typically introduced as multiplications within the attention mechanism. These general rules apply depending on whether encodings are relative or absolute.
To further explore encoding strategies, we introduce an encoding variant that employs concatenation with a linear layer and activation function to revert to the original embedding dimension. To the best of our knowledge, this approach has not been previously utilized in SRS and has seen limited application in NLP. While concatenation benefits include separating semantic and positional information, it introduces a size increase in the architecture, which is particularly critical for Large Language Models. To mitigate this, we incorporate an additional layer, allowing for a gradual introduction of positional information without excessively increasing model complexity.
Concatenated Absolute Encoding
To enhance positional information, we introduce concatenated absolute encoding, which concatenates the positional encoding to the item embedding:
(9) $x=\text{Concat}(x,PE_{(pos,*)})$
A Linear layer is then applied to restore the original embedding dimension without increasing model parameters. This method explicitly incorporates positional information, allowing the model to gradually rely on positional cues and achieving more stable performance (Section 5).
Concatenated Learnable Encoding (Learnt + Con )
Additionally, we apply a concatenation approach to learnable encodings, denoted as concatenated learnable encoding (Learnt + Con ). However, Learnt + Con generally yields worse results compared to the classic residual connection (Section 5), likely because concatenation adds complexity in this case, making it harder for the model to effectively learn embeddings. Furthermore, the gradual introduction of positional information is already inherent in the learnable encoding mechanism.
We do not apply concatenated versions for RoPE and RMHA-4 , as they are relative encodings.
RoPE One Block (RopeOne )
In our experiments (Section 5) we found that RMHA-4 produce more stable results than other encodings. To investigate whether the stability improvements of RMHA-4 were due to introducing positional information at each Transformer block or its role as an RPE -based head, we tested RoPE . The results showed that that RoPE was in a middle point between relative and absolute encodings regarding stability.
We hypothesize that this intermediate stability is due to the application of RoPE across all Transformer blocks (($3$) in Figure 2). To investigate this further, we introduced RopeOne , which applies RoPE exclusively in the first Transformer block (($2$) in Figure 2). As detailed in Section 5, this modification yields slightly improved results. This implies that the rotary aspect, rather than merely being relative, is key to the observed improvements and leads us to the new encoding introduced below.
3.2. Rotatory Encoding
We propose a new encoding method termed Rotatory (Figure 1), with its concatenated variant referred to as Rotatory + Con This approach applies a rotation to define the position of each item. Unlike traditional methods that add positional information during self-attention, our method incorporates these rotations at the initial item embeddings, similar to sinusoidal encoding (Abs ). Vector embeddings offer the advantage of faster training and inference (Press et al., 2022). Therefore, we adopt a formulation similar to RoPE , applying the rotations at the outset. We enhance flexibility by making the rotation angles learnable, allowing the model to capture both absolute and relative positional relationships effectively.
The positional encoding for an item at position $pos$ and embedding dimension $i$ is calculated using a trigonometric approach, modified to include alternating sine signs for a rotational effect. The encoding is defined as follows:
(10) $PE_{(pos,2i)}=(-1)^{i}\sin\left(\theta_{(pos,i)}\right)$
(11) $PE_{(pos,2i+1)}=\cos\left(\theta_{(pos,i)}\right)$
Here, $pos$ is the position of the item in the sequence, $i$ represents the embedding dimension, and $\theta_{(pos,i)}$ is a learnable angle defined as:
(12) $\theta_{(pos,i)}=\frac{E_{pos}}{10000^{(2i/d)}}\cdot 2\pi,\quad\forall pos\in[0,L),\forall i\in[0,H)$
where $E_{pos}\in\mathbb{R}^{L\times H}$ are learnable embeddings. The term $(-1)^{i}$ in the sine function alternates the sign, introducing the rotation effect in the encoding. These positional encodings are then added to the original item embeddings by addition:
(13) $x^{\prime}_{(pos)}=x_{(pos)}+PE_{(pos)}$
where $x_{(pos)}$ is the original semantic embedding of the item at position $pos$, and $PE_{(pos)}$ is the learnable Rotatory  encoding. In the case of Rotatory + Con , the positional encoding is concatenated with the original item embedding and passed through a learned linear transformation for further adaptation. This learnability introduces flexibility, enabling the model to capture both absolute and relative positional information in a dynamic manner, which leads to improved performance in handling complex sequence relationships.
4. Experimental Setup
4.1. Datasets
We selected several Amazon datasets to evaluate performance of various encoding types. For this purpose, we utilised four distinct real-world datasets extracted from product reviews on Amazon.com. These datasets are widely used for SRS under the leave-one-out protocol (He and McAuley, 2016b; Kang and McAuley, 2018b; Hou et al., 2019; Wu et al., 2020; Steck, 2019; Zhou et al., 2020). Since multiple versions of these datasets exist, we opted for the preprocessed versions provided by (Rashed et al., 2022).
Dataset Users Items Interactions Attributes
Men 34,244 110,636 254,870 2,048
Fashion 45,184 166,270 358,003 2,048
Games 31,013 23,715 287,107 506
Beauty 52,204 57,289 394,908 6,507
Table 2. Dataset Statistics.
Dataset Avg Dev Hit Avg Dev NDCG CI-length Density
Men 3.9143 8.2742 5.6686 $6.727\cdot 10^{-5}$
Fashion 4.87 9.7814 9.2257 $4.765\cdot 10^{-5}$
Games 1.124 0.952 1.616 $3.904\cdot 10^{-4}$
Beauty 0.935 0.87 1.22 $1.320\cdot 10^{-4}$
Table 3. Deviations among the different datasets without Encoding. This pattern is generally preserved along the different encodings. The density refers to the inverse of Sparsity.
Covering various product categories, these datasets offer a broad source for training and evaluating recommendation models. They include essential features for each interaction, such as user IDs, item IDs, timestamps, and additional vector-based contextual information. As shown in Tables 2 and 3, they exhibit diversity in terms of sparsity, defined as the ratio of missing user-item interactions. This will be a crucial aspect of our analysis, along with the discreteness of the features.
1.Beauty
(Wu et al., 2020; Zhou et al., 2020; Kang and McAuley, 2018b) The dataset includes discrete and categorical attributes for all beauty products, including fine-grained categories and brands. Mainly categorical and discrete features.
2.Video Games
(Wu et al., 2020; Kang and McAuley, 2018b) The Video Games sub-dataset includes user interactions, reviews, and product details specific to the video game category, such as price, brand and categorical features. Most of the attributes here are discrete and categorical.
3.Men
(He and McAuley, 2016b) The men’s dataset encompasses a comprehensive collection of items falling under men’s clothing. The attributes are dense vectors from image-based features extracted from the last layer of a ResNet50 (Deng et al., 2009) on the ImageNet dataset (He et al., 2016a).
4.Fashion
(He and McAuley, 2016b; Hou et al., 2019) It contains six categories for men’s and women’s clothing. The features were extracted as dense using the same ResNet50 approach.
4.2. Models and Training
To implement CARCA, we converted the original TensorFlow model into PyTorch [^1] https://github.com/researcher1741/Position_encoding_SRS. We used the loss implementation from (Papso, 2023). The original TensorFlow version built upon prior models within the same framework. During our evaluation, we observed that the implementations checked metrics every $20$ epochs, assuming that the optimal performance point would be reached later. This approach was likely adopted due to the extensive number of epochs, exceeding $1000$, which increases the computational cost of evaluations. To achieve more reliable evaluations, we implemented a logarithmic measurement of the metrics, which resulted in slightly improved outcomes in certain cases.
As in previous works (Rashed et al., 2022; Kang and McAuley, 2018a), we use an ADAM optimiser to minimise the binary cross-entropy loss of the CARCA model while masking the padded items to prevent them from contributing to the loss function. For a given sequence of items interactions for a user $u$ as $\{i^{u}_{1},\ldots,i^{u}_{N}\}$, we create an input list as $I^{u+}=\{i^{u}_{1},\ldots,i^{u}_{N-1}\}$ by deleting the last item of the list, a positive target list as $T^{u+}=\{i^{u}_{2},\ldots,i^{u}_{N}\}$ by shifting the input list by one, and a negative target list as $T^{u-}=\{i^{rand_{1}},\ldots,i^{rand_{N-1}}\}$ generated by random negative items. Then the loss is given by:
Encoding Avg Dev Hit Avg Dev NDCG runs CI-length
Abs 5.43 8.25 9.38 6.71
Abs + Con 3.52 5.56 7.12 5.35
Learnt 4.79 7.64 7.88 7.05
Learnt + Con 3.17 6.36 6.62 4.42
None 2.35 5.26 8.00 3.29
RMHA-4 0.51 1.08 6.00 0.88
RoPE 2.38 5.45 5.00 4.04
Rotatory 2.46 5.01 8.38 3.48
Rotatory + Con 2.48 5.18 6.75 3.49
RopeOne 2.82 6.54 4.29 5.46
RoPE-Longer 3.33 8.15 5.25 5.99
Rotatory-Longer 1.05 1.04 4.67 2.08
Rotatory + Con-Longer 0.99 1.24 6.00 1.59
RMHA-4-Longer 0.71 4.00 6.00 1.14
Table 4. Deviations among the different encodings for the default upper bound values ($0.0001$ for all except NaN for Games). Relative encoding in all the heads, RMHA-4 , presents more stability, followed by our Rotatory versions.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
@@ 4.2. Models and Training
Beauty silu Rotatory + Con 0.0001 61.68 1.04 42.31 0.97
Beauty leaky Rotatory + Con 0.0001 61.16 0.77 42.83 0.48
@@
Games leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93
@@
Table 5. The table displays the top $4$ results for each dataset on cases with a standard deviation for Hit Meanbelow $3$.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
@@
Games leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93
@@
Table 6. The table displays the top $4$ results for each dataset on cases with a standard deviation for Hit Meanbelow $12$.
4.3. Metrics
We employ two widely used metrics for Next-Item Recommendation and SRS: Hit@10 and Normalised Discounted Cumulative Gain (NDCG).
Hit@10: This metric measures the fraction of times the ground truth next item appears within the top $10$ recommended items. For simplicity, we refer to Hit@10 as Hit, its mean as Hit Mean, its devisation as Hit Dev  and the average of the deviations as Avg Dev Hit.